# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v1

@@ Acknowledgements
We thank Kevin Zielnicki, Molly Davies, Sven Schmit, Brad Klingenberg, Chris Rinaldi, Viridiana Lourdes, Izzy Farley, Sudhish Kasaba Ramesh, Sean Kelly, and Adith Swaminathan for helpful discussions and feedback on earlier drafts of this work.
@@ Appendix B Algorithms (extended)
This appendix gives compact, cross-fitted pseudocode for CJE modules: reward calibration (AutoCal-R), surrogate-indexed weight calibration (SIMCal-W), estimators (Cal-IPS, DR-CPO), IF-space stacking, and oracle–uncertainty–aware variance (OUA). We reuse the same $K$-fold map $F(i)\!\in\!\{1{:}K\}$ across all modules. “OOF” denotes out-of-fold predictions used along the IF path.
@@
PAVA is $O(n)$ after a shared sort by $S$. SIMCal-W is linear-time per fold; the stacking QP is $3{\times}3$ (weights) or a small $|\mathcal{E}|{\times}|\mathcal{E}|$ system. DR-CPO adds one rollout + judge per $(X,\pi^{\prime})$. The OUA jackknife refits the calibrator $K$ times and reruns the pipeline; TF caches and precomputed features amortize cost.
@@ C.2 Isotonic regression: mean preservation and majorization
After sorting by $s$, the isotonic fitted vector $\hat{u}$ is a mean-preserving adjacent pooling of $y$; hence for any convex $\phi$, $\sum_{i}\phi(\hat{u}_{i})\leq\sum_{i}\phi(y_{i})$ (HardyLittlewoodPolya1952; MarshallOlkinArnold2011). In particular, with sample mean one, $\operatorname{Var}_{n}(\hat{u})\leq\operatorname{Var}_{n}(y)$ and $\mathrm{ESS}(\hat{u})\geq\mathrm{ESS}(y)$.
Proofs are standard; see Ayer1955PAVA; Barlow1972; RobertsonWrightDykstra1988; Banerjee2001.
@@ C.3 Proof of Theorem 1 (surrogate EIF & variance drop)
with $q^{\star}_{R}(x,a)=\mathbb{E}[R^{\star}\mid x,a]$ and $g^{\star}_{\pi^{\prime},R}(x)=\mathbb{E}_{A\sim\pi^{\prime}(\cdot\mid x)}[q^{\star}_{R}(x,A)]$ (Bickel1993; VaartWellner2000). Since $m^{\star}$ is the $L^{2}$ projection of $W_{\pi^{\prime}}$ onto $L^{2}(\sigma(S))$, Pythagoras (or (3)) implies $\operatorname{Var}(\phi_{\mathrm{sur}})\leq\operatorname{Var}(\phi_{\mathrm{uncon}})$, strictly unless $W_{\pi^{\prime}}\in L^{2}(\sigma(S))$ and $R^{\star}$ is degenerate.
@@ C.4 Proof of Theorem 2 (Efficiency via Model Restriction)
For attainability, standard semiparametric arguments apply: one-step corrections or TMLE with cross-fitting yield regular estimators achieving the restricted efficiency bound (Bickel1993; VaartWellner2000; vanDerLaanRose2011TL).
@@ C.11 Auxiliary lemmas used in the main proofs
Metric projections onto closed convex sets in Hilbert spaces are firmly non-expansive: $\|\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y)\|^{2}\leq\langle\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y),\,x-y\rangle$. Hence compositions of DbP modules (reward, weight, IF-space projections) are non-expansive. See, e.g., BauschkeCombettes2017.
@@ C.12 Coverage-Limited Efficiency (CLE) Diagnostic
C.13 Assumptions ledger
@@
The calibration function $f$ transports across environments $g\in\mathcal{G}$: $\mathbb{E}_{g}[Y\mid X,A,S]=f(S,X)$ for all $g$ (PearlBareinboim2014).
@@
C.14 What the bounds do not include
@@
C.15 Citations for technical ingredients
Semiparametric efficiency and one-step/TMLE: Bickel1993; VaartWellner2000; vanDerLaanRose2011TL; Tsiatis2006.
Isotonic regression / PAVA and order-restricted inference: Ayer1955PAVA; Barlow1972; RobertsonWrightDykstra1988; Banerjee2001.
Majorization theory: HardyLittlewoodPolya1952; MarshallOlkinArnold2011.
Calibration for OPE: KallusMao2020Surrogates; vanDerLaanLinCaroneLuedtke2025StabIPW; vanDerLaanLuedtkeCarone2024DRCalibration.
Jackknife/bootstraps for dependence: Kunsch1989BlockBootstrap; PolitisRomano1994StationaryBootstrap.
Projections and non-expansive maps: BauschkeCombettes2017.
@@ D.4 Uncertainty: IF variance and OUA addition
When time/cluster dependence is suspected, also report: (i) cluster-robust sandwich SEs when a cluster id (e.g., session/user) is available; and (ii) block/stationary bootstrap intervals (block length chosen by a simple variance-stability sweep) (Kunsch1989BlockBootstrap; PolitisRomano1994StationaryBootstrap).
@@ E.9 Exact prompts used in experiments
Both the judge (GPT-4.1-nano) and oracle (GPT-5) use identical prompts; the only difference is the underlying model. Temperature = 0.0 for judge, 1.0 for oracle (model constraint).
@@ Appendix F Efficient Budget Allocation
Using costs from our experiments: the GPT-4.1-nano judge is approximately $16\times$ cheaper than the GPT-5 oracle, giving a cost ratio $c_{S}/c_{Y}\approx 0.064$. At $n=1000$ prompts with $m=50$ oracle labels (5% oracle fraction), the observed OUA share is $\omega\approx 55\%$.

### v2

@@ Acknowledgements
We thank Kevin Zielnicki, Molly Davies, Sven Schmit, Brad Klingenberg, Chris Rinaldi, Viridiana Lourdes, Izzy Farley, Sudhish Kasaba Ramesh, Sean Kelly, Adith Swaminathan, Jasmine Nettiksimmons, Winston Chou, Manjari Narayan, and Tom Cunningham for helpful discussions and feedback on earlier drafts of this work.
@@ Appendix B Algorithms (extended)
This appendix gives compact, cross-fitted pseudocode for CJE modules: reward calibration (AutoCal-R), surrogate-indexed weight calibration (SIMCal-W), estimators (Cal-IPS, DR-CPO), IF-space stacking, oracle–uncertainty–aware variance (OUA), and cluster bootstrap for Direct mode (Bootstrap-Direct). We reuse the same $K$-fold map $F(i)\!\in\!\{1{:}K\}$ across all modules. “OOF” denotes out-of-fold predictions used along the IF path.
@@
Algorithm 7 Bootstrap-Direct: cluster bootstrap with calibrator refit for Direct mode
1:Inputs: Eval data $\{(X_{i},S_{i},Y_{i},C_{i})\}_{i=1}^{n}$ with cluster IDs $C_{i}\in\{1,\ldots,M\}$; calibrator mode $m$ (fixed from full-data selection); oracle mask $L$; replicates B=2000B{=}2000; min oracle count $n_{\min}{=}30$
2:Outputs: Point estimates $\{\hat{\theta}_{p}\}$, 95% CIs for each policy $p$
3: Fit calibrator $\hat{f}$ on oracle slice; compute $\hat{\theta}_{p}^{\mathrm{aug}}$ for each policy via Equation 3
4:for $b=1$ to $B$ do
5:  repeat
6:   Draw $M$ clusters with replacement: $\{C_{1}^{*},\ldots,C_{M}^{*}\}$
7:   Form bootstrap sample $\mathcal{D}^{*}=\bigcup_{j=1}^{M}\{i:C_{i}=C_{j}^{*}\}$
8:  until $|L\cap\mathcal{D}^{*}|\geq n_{\min}$ {Resample-until-valid}
9:  Refit calibrator $\hat{f}^{(b)}$ on $\{(S_{i},Y_{i}):i\in L\cap\mathcal{D}^{*}\}$ with mode $m$
10:  Compute $\hat{\theta}_{p}^{(b)}$ for each policy using $\hat{f}^{(b)}$ and $\hat{\theta}_{\mathrm{aug}}$ on $\mathcal{D}^{*}$
11:end for
12:Percentile CIs: $\hat{\theta}_{p}^{\mathrm{lo}},\hat{\theta}_{p}^{\mathrm{hi}}]=[\hat{\theta}_{p}^{(0.025B)},\hat{\theta}_{p}^{(0.975B)}]$
@@
PAVA is $O(n)$ after a shared sort by $S$. SIMCal-W is linear-time per fold; the stacking QP is $3{\times}3$ (weights) or a small $|\mathcal{E}|{\times}|\mathcal{E}|$ system. DR-CPO adds one rollout + judge per $(X,\pi^{\prime})$. The OUA jackknife refits the calibrator $K$ times and reruns the pipeline. Bootstrap-Direct requires B=2, [^0] B{=}2{,}000 refits of AutoCal-R; since isotonic regression is $O(n\log n)$, total cost is $O(Bn\log n)$. TF caches and precomputed features amortize cost.
@@ C.2 Isotonic regression: mean preservation and majorization
After sorting by $s$, the isotonic fitted vector $\hat{u}$ is a mean-preserving adjacent pooling of $y$; hence for any convex $\phi$, $\sum_{i}\phi(\hat{u}_{i})\leq\sum_{i}\phi(y_{i})$ (Hardy et al., 1952; Marshall et al., 2011). In particular, with sample mean one, $\operatorname{Var}_{n}(\hat{u})\leq\operatorname{Var}_{n}(y)$ and $\mathrm{ESS}(\hat{u})\geq\mathrm{ESS}(y)$.
Proofs are standard; see Ayer et al. (1955); Barlow et al. (1972); Robertson et al. (1988); Banerjee (2001).
@@ C.3 Proof of Theorem 1 (surrogate EIF & variance drop)
with $q^{\star}_{R}(x,a)=\mathbb{E}[R^{\star}\mid x,a]$ and $g^{\star}_{\pi^{\prime},R}(x)=\mathbb{E}_{A\sim\pi^{\prime}(\cdot\mid x)}[q^{\star}_{R}(x,A)]$ (Bickel et al., 1993; van der Vaart & Wellner, 2000). Since $m^{\star}$ is the $L^{2}$ projection of $W_{\pi^{\prime}}$ onto $L^{2}(\sigma(S))$, Pythagoras (or (5)) implies $\operatorname{Var}(\phi_{\mathrm{sur}})\leq\operatorname{Var}(\phi_{\mathrm{uncon}})$, strictly unless $W_{\pi^{\prime}}\in L^{2}(\sigma(S))$ and $R^{\star}$ is degenerate.
@@ C.4 Proof of Theorem 2 (Efficiency via Model Restriction)
For attainability, standard semiparametric arguments apply: one-step corrections or TMLE with cross-fitting yield regular estimators achieving the restricted efficiency bound (Bickel et al., 1993; van der Vaart & Wellner, 2000; van der Laan & Rose, 2011).
@@ C.11 Auxiliary lemmas used in the main proofs
Metric projections onto closed convex sets in Hilbert spaces are firmly non-expansive: $\|\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y)\|^{2}\leq\langle\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y),\,x-y\rangle$. Hence compositions of DbP modules (reward, weight, IF-space projections) are non-expansive. See, e.g., Bauschke & Combettes (2017, Prop. 4.16).
@@ C.12 Coverage-Limited Efficiency (CLE) Diagnostic
C.13 EIF derivation for Direct mode
Direct mode estimates $\theta:=\mathbb{E}[Y]$ when oracle labels are observed only for a subset $L\subset[n]$. Let $L_{i}\in\{0,1\}$ indicate oracle observation, and assume missing at random (MAR): $\Pr(L_{i}=1\mid Z_{i},Y_{i})=\Pr(L_{i}=1\mid Z_{i})=:e(Z_{i})$, where $Z_{i}$ is the calibration index. Under this model, the efficient influence function for $\theta$ is (Tsiatis, 2006; Bang & Robins, 2005):
$\phi(O_{i})=\mu(Z_{i})-\theta+\frac{L_{i}}{e(Z_{i})}\bigl(Y_{i}-\mu(Z_{i})\bigr),$
[…]
$\hat{\theta}_{\mathrm{aug}}=\frac{1}{n}\sum_{i=1}^{n}\hat{\mu}(Z_{i})+\frac{1}{|L|}\sum_{i\in L}\bigl(Y_{i}-\hat{\mu}^{(-i)}(Z_{i})\bigr),$
which matches Equation 3. The cross-fit predictions $\hat{\mu}^{(-i)}$ avoid overfitting bias in the residual term.
Remark (TMLE equivalence). The cross-fitted augmented estimator $\hat{\theta}_{\mathrm{aug}}$ is the one-step/AIPW estimator for the EIF above, specializing to $e(Z)\approx|L|/n$ under approximately uniform oracle sampling. A TMLE that targets $\hat{\mu}$ via a logistic fluctuation using clever covariate $H(O)=L/e(Z)$ (equivalently, $H(Z)=1/e(Z)$ on labeled rows) solves the EIF estimating equation and is asymptotically equivalent to $\hat{\theta}_{\mathrm{aug}}$, yielding the same influence function and achieving the semiparametric efficiency bound for $\theta$ in the MAR model (van der Laan & Rose, 2011).
C.14 Assumptions ledger
@@
The calibration function $f$ transports across environments $g\in\mathcal{G}$: $\mathbb{E}_{g}[Y\mid X,A,S]=f(S,X)$ for all $g$ (Pearl & Bareinboim, 2014).
@@
C.15 What the bounds do not include
@@
C.16 Citations for technical ingredients
Semiparametric efficiency and one-step/TMLE: Bickel et al. (1993); van der Vaart & Wellner (2000); van der Laan & Rose (2011); Tsiatis (2006).
Isotonic regression / PAVA and order-restricted inference: Ayer et al. (1955); Barlow et al. (1972); Robertson et al. (1988); Banerjee (2001).
Majorization theory: Hardy et al. (1952); Marshall et al. (2011).
Calibration for OPE: Kallus & Mao (2020); van der Laan et al. (2025; 2024).
Jackknife/bootstraps for dependence: Künsch (1989); Politis & Romano (1994).
Projections and non-expansive maps: Bauschke & Combettes (2017).
@@ D.4 Uncertainty: IF variance and OUA addition
When time/cluster dependence is suspected, also report: (i) cluster-robust sandwich SEs when a cluster id (e.g., session/user) is available; and (ii) block/stationary bootstrap intervals (block length chosen by a simple variance-stability sweep) (Künsch, 1989; Politis & Romano, 1994).
Bootstrap with bias-corrected estimation.
For Direct mode, we recommend bootstrap inference with $\hat{\theta}_{\mathrm{aug}}$ (see Section 3.5). Table 7 compares UQ methods across oracle fractions, showing that bootstrap with $\hat{\theta}_{\mathrm{aug}}$ achieves $\sim$}95% coverage by correcting both variance and bias.
[…]
Oracle Method Coverage $\bar{Z}$ SD(Z)(Z)
5% Cluster-Robust 22% +1.52+1.52 9.01
OUA Jackknife 80% +0.48+0.48 3.11
Bootstrap + $\hat{\theta}_{\mathrm{aug}}$ 95% +0.10+0.10 0.98
10% Cluster-Robust 30% +1.12+1.12 6.34
OUA Jackknife 80% +0.40+0.40 2.52
Bootstrap + $\hat{\theta}_{\mathrm{aug}}$ 96% +0.15+0.15 0.98
25% Cluster-Robust 42% +0.36+0.36 3.46
OUA Jackknife 84% +0.18+0.18 2.09
Bootstrap + $\hat{\theta}_{\mathrm{aug}}$ 97% +0.08+0.08 0.95
50% Cluster-Robust 56% 0.07-0.07 2.63
OUA Jackknife 87% 0.05-0.05 1.96
Bootstrap + $\hat{\theta}_{\mathrm{aug}}$ 97% +0.11+0.11 0.92
$Z=(\hat{\theta}-\theta)/\widehat{\mathrm{SE}}$ is the normalized error. N=600N{=}600 experiments per cell (4 sample sizes $\times$ 50 seeds $\times$ 3 policies). $\hat{\theta}_{\mathrm{aug}}$ corrects the plug-in bias via AIPW-style residual augmentation.
@@ E.9 Exact prompts used in experiments
Both the judge (gpt-4.1-nano-2025-04-14) and oracle (gpt-5-2025-08-07) use identical prompts; the only difference is the underlying model. Temperature = 0.0 for judge, 1.0 for oracle (temperature 0 was not available for gpt-5-2025-08-07 at experiment time).
@@ Appendix F Efficient Budget Allocation
Using costs from our experiments: the gpt-4.1-nano-2025-04-14 judge is approximately $16\times$ cheaper than the gpt-5-2025-08-07 oracle, giving a cost ratio $c_{S}/c_{Y}\approx 0.064$. At $n=1000$ prompts with $m=50$ oracle labels (5% oracle fraction), the observed OUA share is $\omega\approx 55\%$.